OpenStudy (anonymous):

solve by completing the square x^2 - 14x +3 = 0

OpenStudy (anonymous):

now this makes sense \[x^2-14x+3=0\] \[x^2-14x=-3\] \[(x-7)^2=-3+7^2\] \[(x-7)^2=46\] \[x-7=\pm\sqrt{46}\] \[x=7\pm\sqrt{46}\]